3. In the following table, fill in the blanks after you have completed the entire table, determine the profit-maximizing output. Feel free to create columns in excel and provide the answer.

Output Price Total Revenue Marginal Revenue Total Cost Marginal Cost Total Profit
1 $20 $40 ($20)
2 50
3 60
4 65 $5
5 85
6 $120 120
Answer #1
Output(Q) Price TR($)= P*Q MR($) = ∆TR/∆Q TC($) MC($) = ∆TC/∆Q Total Profit($) = TR - TC
1 20 20 40 -20
2 20 40 20 50 10 -10
3 20 60 20 60 10 0
4 20 80 20 65 5 15
5 20 100 20 85 20 15
6 20 120 20 120 35 0